OpenStudy (anonymous):

The expression (2x2)3 is equivalent to _____. Note- The "2" after the variable is an exponent and the "3" outside of the parentheses is also an exponent.

hartnn (hartnn):

so, \((2x^2)^3\) right ?? first use, \((mn)^a = m^an^a\) so what about \((2x^2)^3\) ? treat m as 2 and n as x^2

OpenStudy (anonymous):

\[6^{2} or 6^{5} or 8^{6}\]

hartnn (hartnn):

ohh,, there is no x there! right ?

OpenStudy (anonymous):

Yes there is. I accidentally forgot to put the x's in the choices :/

hartnn (hartnn):

lets first see the constant part , since you have choices the exponent will get distributed to the 2 and x^2 term both so, whats 2^3 =.. ?

OpenStudy (anonymous):

\[6x ^{2}\] or \[6x ^{5}\] or\[8x ^{6}\]

hartnn (hartnn):

whats 2^3 = ...

hartnn (hartnn):

:O 2^3 = 2*2*2 = ... not 6

OpenStudy (anonymous):

Lol my bad. It's 8

hartnn (hartnn):

yeah and just to make sure \((a^b)^c=a^{bc}\) so, \((x^2)^3 = x^6\) so its 8x^6